OpenStudy (anonymous):

add simplify if possible (2 over 25x^3)+(1 over 10z^2)

OpenStudy (anonymous):

\[\frac{1}{50} \left(\frac{4}{x^3}+\frac{5}{z^2}\right) \text{ or } \frac{5 x^3+4 z^2}{50 x^3 z^2} \text{ together} \]

jhonyy9 (jhonyy9):

2/25x*3 + 1/10z*2 =(4z*2 +5x*3)/50x*(3)z*(2)

OpenStudy (anonymous):

@jhonyy9 Shouldn't your expression be written as follows? 2/25x^3 + 1/10z^2 =(4z^2 +5x^3)/50x^(3)z^(2)